OpenStudy (anonymous):

Which of the following represents the remainder when dividing f(x) = xsqrt3 - 8x -10 by the factor x-2 ? A. −18 B. −12 C. −4 D. −2

OpenStudy (anonymous):

a. -18 b. -12 c. -4 d.-2

OpenStudy (anonymous):

@waterineyes you know this type of question

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

\(f(x) = x \sqrt{3} - 8x - 10\) ??

OpenStudy (anonymous):

NO, I think it is : \(f(x) = x^3 - 8x - 10\) .. :)

OpenStudy (anonymous):

oh srry my internet was acting weird

OpenStudy (anonymous):

so you plug in the answer choices @waterineyes

OpenStudy (anonymous):

What is the f(x) value?? I have written two values, what is real one?

OpenStudy (anonymous):

is it -2

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

(x-2) is a factor means \(x = 2\), so you have to put \(x=2\) in that equation, and find out what you are left with.